OpenStudy (anonymous):

simplify by expressing fractional exponents instead of radical 8root of x^2 y^4

OpenStudy (anonymous):

\[8\sqrt{x^2y^4}=8(x^2y^4)^{\frac{1}{2}}\]

OpenStudy (anonymous):

Just use properties of exponents: \[(a^b)^c=a^{bc}\]

OpenStudy (anonymous):

I tried it anf I still don`t get it, it gave me 8(abs(^nee) sq. rt of y^4 I`m lost

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

yeaah D:

OpenStudy (anonymous):

do you understand \(\large (a^b)^c = a^{bc} \) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok... so, \(\huge 8\sqrt{x^2y^4}=8(x^2y^4)^\frac{1}{2}=8(x^2)^\frac{1}{2} (y^4)^\frac{1}{2} \) \(\huge 8(x^{2\times\frac{1}{2}})(y^{4\times\frac{1}{2}}) \)

OpenStudy (anonymous):

can you simplify it from here?

OpenStudy (anonymous):

I`m not sure... or the first one I get 8 and for the second one I get y^2, would it be 8y^2?

OpenStudy (anonymous):

*for

OpenStudy (anonymous):

umm.... actually, this is what I should have shown..